### v1

@@ 3.3 Error Analysis and Quantizer Design for Model Learning and Quantized Q-Learning
As noted above, the weighting measure depends on the exploration policy, quantizer and system model, and cannot be assigned arbitrarily. This dependence introduces a key difficulty in bounding the expected loss, as the weighing measure over the overflow bin, i.e. $\pi_{y_{M+1}}$ can no longer be freely chosen or controlled. This limitation is significant for the analysis of non-compact spaces. Nevertheless, we will show that it is still possible to establish a meaningful upper bound on the expected error by employing a Lyapunov-based analysis, analogous to the approach presented in Section 2.2.
We first start with the following result which relates the approximation error to occupation measures.
In the previous section, we studied the approximate model where the weighting measures $\hat{\pi}^{*}_{y_{i}}$ for each quantization bin were chosen freely and $\hat{\pi}^{*}_{y_{i}}$ were chosen as Dirac measures concentrated at the $\ell_{1}$ centroids (medians) of each bin $B_{i}$. However, when implementing Q-learning in the POMDP framework as described in this section, we lose the freedom to choose these weighting measures independently. The measures $\hat{\pi}^{*}_{y_{i}}$’s are inherently determined by the invariant distribution $\pi_{\gamma^{*}}$ under the exploration policy $\gamma^{*}$. We note that under weak continuity and compactness of the state space conditions, the set of invariant measures is compact in the weak topology. By considering this fact, we introduce the following theorem:
@@
Under Assumption Assumption 2.2 and Assumption 3.1, given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}\in\mathcal{X}$:
$\left|J^{*}_{\beta}(x_{0})-\min_{v}\{Q^{*}(x_{0},v)\}\right|$
$\quad\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\int_{B_{i}}\|x-x^{\prime}\|\mu_{\beta}(dx)\hat{\pi}_{y_{i}}(dx^{\prime}),$ (24)
@@
Assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
@@
where $V(x)=|x|^{m}$, $m>1$. Then, under Assumption 2.1 and Assumption 2.2, provided that the cost function $c$ is bounded, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathbb{X}$:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1-1/m})(1-\beta)}\right),$ (27)
@@
$C:=\frac{|x_{0}|^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
@@ 4 Concluding Remarks
For MDPs with unbounded state spaces we presented upper bounds on finite model approximation errors via optimizing the quantizers used for finite model approximations. We also considered implications on quantized Q-learning and the performance of policies obtained via Q-learning where the quantized state is treated as the state itself. We noted the distinctions between planning, where approximating MDPs can be independently designed, and learning, where approximating MDPs are restricted to be defined by invariant measures of Markov chains under exploration policies, leading to significant subtleties on quantizer design performance. Nonetheless, asymptotic near optimality can be established under both setups with explicit convergence rates. In particular, under Lyapunov growth conditions, we obtained explicit upper bounds which decay to zero as the number of bins approaches infinity.
@@ Appendix C Proof of Theorem 7
Step 1: Let $\mathds{X}\subseteq\mathbb{R}$ be the state space, which we partition into $M+1$ quantization bins $\{B_{1},B_{2},\ldots,B_{M},B_{M+1}\}$. The first $M$ bins cover a compact subset $\mathcal{K}=[-\frac{N}{2},\frac{N}{2}]\subset\mathds{X}$, and the last bin $B_{M+1}$ is an overflow bin that captures the rest of the state space outside $\mathcal{K}$.
We apply a uniform quantizer to the compact region $\mathcal{K}$, dividing it into $M$ bins of equal length. The quantization width of each bin is:
@@
where $\mu_{\beta}$ is the normalized discounted occupation measure under a stationary policy $\gamma_{s}$.
Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
@@
Step 3: In the overflow bin $B_{M+1}$, the state space may be unbounded. In what follows, to simplify the analysis, we will assume that the overflow bin is always mapped to state $x=0$. We need to bound:
@@
where $\hat{\pi}^{*}_{M+1}(dx^{\prime})$ is the normalized weighing measure over the bin $B_{M+1}$ and where the last step follows since we map the overflow bin directly to 0, i.e. $x^{\prime}=1$ with probability 1.
Using Hölder’s inequality to bound the expected loss over $B_{M+1}$:
@@
Step 4: In the following, we bound the moment under the measure $\mu_{\beta}^{\gamma_{s}}$. Define the process $\{M_{t}\}$ for $t\geq 0$:
@@
with $M_{0}=V(x_{0})-b=|x_{0}|^{m}-b.$
Observe the following with respect to the filtration $\{\mathcal{F}_{t}\}$, where $\mathcal{F}_{t}=\sigma(X_{1},X_{2},\dots,X_{t})$ is the natural filtration:
@@
$\leq\frac{(1-\alpha)V(x_{t})}{(1-\alpha)^{t+1}}-\frac{b}{(1-\alpha)^{t+1}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\frac{b}{(1-\alpha)^{t+1}}$
$=0,$
@@
$=(V(X_{0})-b)(1-\alpha)^{t}+b\sum_{k=0}^{t}(1-\alpha)^{k}$
$=(V(X_{0})-b)(1-\alpha)^{t}+b\cdot\frac{1-(1-\alpha)^{t+1}}{\alpha}.$
@@
which shows that the error bound $C$ depends on the initial state $x_{0}$ and other constants.
@@
where $C$ is defined in (35).
Combining the bounds from (32), (C) and (36), the total expected loss is bounded by:
@@
Appendix D Proof of Theorem 8
Proof.
Step 1: Let the state space be $\mathds{X}\subseteq\mathbb{R}^{n}$ and fix a side–length parameter $N>0$. Choose an integer $k\geq 1$ and set
@@ Appendix D Proof of Theorem 8
which is the set of states outside the compact (granular) grid.
Step 2: For any state $x\in\mathcal{K}$, due to the uniform quantization, we have that
$L(x)\leq{n\Delta}$
We decompose the expected loss into two parts as in the proof of Theorem 7:
@@
Step 3 : Inside the overflow bin $B_{M+1}\subseteq\mathds{X}$ the state space is unbounded. In the quantizer, for simplicity in analysis, we may take that the overflow bin is directly mapped to state $x=0$. We must bound:
@@
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\bigl(\|x\|_{1}+\|x^{\prime}\|_{1}\bigr)\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]$
@@
Step 4: Repeating the arguments in Step 4 of the proof of Theorem 7, we can show that
@@
We select $N\;=\;2\bigl(C\,M\bigr)^{1/m}.$ By Markov’s Inequality:
@@
$\leq\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]}{(\tfrac{N}{2})^{m}}=\frac{C}{C\,M}=\frac{1}{M}.$
@@
$\Delta=\frac{N}{k}=2C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}.$
Inserting these bounds into (43) and assuming that $m>n$:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}\left(\frac{1}{M}\right)^{1-\frac{1}{m}}$
$\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}M^{-\left(1-\frac{1}{m}\right)}$
$\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}$
$=(2n+1)C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}$
$=(2n+1)C^{1/m}M^{-\left(\frac{1}{n}-\frac{1}{m}\right)}.$
As a result, as $M\to\infty$, the total expected loss $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\to 0$ since $m>n$.
Under necessary conditions, we obtain:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{1/n-1/m})(1-\beta)},$ (44)
where $n$ is the dimension of the state space, $M$ is the number of quantization bins, $m>n$ is the moment in the Lyapunov function, and $C$ is (defined in (35)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$. ∎
Appendix E Proof of Theorem 9
@@
Letting $\alpha:=(1-\mu(\mathds{X}))$, and repeating this step, we write
$\|h_{k}\|_{\infty}\leq\|c\|_{\infty}\sum_{t=0}^{k-1}\alpha^{t}.$
Finally, using the fact that the operator $\hat{T}$ is a contraction under the supremum norm under Assumption 2.3 with the fixed point $\hat{h}$, we conclude
$\|\hat{h}\|_{\infty}\leq\frac{\|c\|_{\infty}}{1-\alpha}=\frac{\|c\|_{\infty}}{\mu(\mathds{X})}.$
Identical steps can be used to show that $\|h\|_{\infty}\leq\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$. Combining what we have so far, we write
[…]
$\leq\left(\alpha_{c}+\frac{\|c\|_{\infty}\alpha_{T}}{\mu(\mathds{X})}\right)\int L(x)\pi_{\gamma_{s}}(dx)$
which completes the proof.
@@ Appendix F Proof of Theorem 11
First we assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
$\sup_{x\in\mathds{X},u\in\mathds{U}}\mathbb{E}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$ (48)
where $f(x)=|x|^{m}$, $m>1$.
Then, we choose a collection of quantization bin $\{B_{i}\}_{i}^{M+1}$ such that the first $M$ bins quantize the compact set $(bM)^{1/m},(bM)^{1/m}]$ uniformly and the last bin $B_{M+1}$ is the overflow bin that captures the rest of the state space. Then, the total expected distortion becomes:
In the overflow bin $B_{M+1}$, the state space may be unbounded, and the quantization error $L(x)=\|x-y_{M+1}\|$ can be large. We need to bound:
$\int_{B_{M+1}}L(x)\,\pi_{\gamma_{s}}(dx)=\int_{B_{M+1}}|x-y_{M+1}|\,\pi_{\gamma_{s}}(dx),$ (49)
where $y_{M}$ is the representative point for bin $B_{M+1}$.
We choose $y_{M+1}=0$ (or any fixed point), so that $L(x)=|x|$ for $x\in B_{M+1}$.
By our assumption, there exists a Lyapunov function $V$ that satisfies inequality (48). Then, by Theorem 10, for any initial state $x_{0}=x$, we have:
$\mathbb{E}_{x}\left[\sum_{t=0}^{T-1}f(x_{t})\right]\leq V(x)+\mathbb{E}\left[\sum_{t=0}^{T-1}b\right]=V(x)+bT,$
where $T$ is a deterministic stopping time. By taking the limit for both sides, we get:
$\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{x}\left[\sum_{t=0}^{T-1}f(x_{t})\right]\leq\limsup_{T\to\infty}\frac{1}{T}(V(x)+bT)=b,$
Now, consider any invariant probability measure $\pi$. Recall that we assigned $f=|x|^{m},m>1$. Fix $N>0$ and define $f_{N}(x)=\min(|x|^{m},N)$. Then, we can write (see [34, p. 75]):
$\int_{\mathds{X}}f_{N}(x)\pi(dx)$ $=\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{x_{0}\sim\pi}\left[E_{x_{0}}[f_{N}(X_{t})]\right]$
$\leq\mathbb{E}_{\pi}\left[\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}E_{x_{0}}[f_{N}(X_{t})]\right]\leq b,$
where the first inequality comes from applying Fatou’s Lemma since $f_{N}(x)$ is bounded by $N$. By Taking the limit $N\to\infty$ and noting that $f_{N}(x)\to|x|^{m}$ pointwise and monotonically, by Monotone Convergence Theorem, we obtain $\int_{\mathds{X}}|x|^{m}\pi(dx)\leq b$.
Finally, since this holds for any invariant probability measure $\pi$ and $|x|^{m}$ is non-negative, we conclude that $\mathbb{E}[|X_{t}|^{m}]\leq b,\forall t\geq 0.$ By following the same procedure in the proof of Theorem 7, we obtain the following bound:
$\leq\Delta\ +\frac{b^{1/m}}{M^{1-1/m}}\leq\frac{2b^{1/m}}{M^{1-1/m}}+\frac{b^{1/m}}{M^{1-1/m}}=\frac{3b^{1/m}}{M^{1-1/m}}.$ (51)
$\left|\hat{J}_{avg}(x_{0})-J^{*}_{avg}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{3b^{1/m}}{(M^{1-1/m})},$ (52)
Appendix G Proof of Theorem 14
Proof.
Step1: We partition the state space similar to the quantization in Theorem 7 by choosing the compact set $\mathcal{K}=[-(CM)^{1/m},(CM)^{1/m}]$. Then, for any state $x\in\mathcal{K}$, the quantization error $L(x)$ satisfies:
$L(x)\leq{\Delta}=\frac{2(CM)^{1/m}}{M}=\frac{2C^{1/m}}{M^{1-1/m}}.$ (53)
@@ Appendix G Proof of Theorem 14
$\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}(\mathcal{K})=\frac{2C^{1/m}}{M^{1-1/m}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (55)
@@
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}|x-x^{\prime}|\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (56)
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\left(|x|+|x^{\prime}|\right)\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (57)
$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]+\mathbb{E}_{\hat{\pi}_{M+1}}[|X^{\prime}|]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1}),$ (58)
@@
Observe that if $x_{0}\sim\pi_{\gamma^{*}}$ then $\mu_{\beta}^{\gamma_{s}}=\pi_{\gamma^{*}}$ and therefore the terms in the summation above, that is,
$\mathbb{E}_{\hat{\pi}^{*}_{M+1}}[|X^{\prime}|]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mu_{\beta}^{\gamma_{s}}(B_{M+1})\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]$
will be identical.
@@
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\mathbb{E}[|X|\mathds{1}_{B_{M+1}}(X)|X_{0}=x]]\leq\hat{C}^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$
@@
$\hat{C}:=\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|x_{0}|]^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$

### v2

@@ 3.3 Error Analysis and Quantizer Design for Model Learning and Quantized Q-Learning
As noted above, the weighting measure depends on the exploration policy, quantizer and system model, and cannot be assigned arbitrarily. This dependence introduces a key difficulty in bounding the expected loss, as the weighing measure over the overflow bin, i.e. $\pi_{y_{M+1}}$ can no longer be freely chosen or controlled. This limitation is significant for the analysis of non-compact spaces.
In the previous section, we studied the approximate model where the weighting measures $\hat{\pi}^{*}_{y_{i}}$ for each quantization bin were chosen freely and $\hat{\pi}^{*}_{y_{i}}$ were chosen as Dirac measures concentrated at the $\ell_{1}$ centroids (medians) of each bin $B_{i}$. However, when implementing Q-learning in the POMDP framework as described in this section, we lose the freedom to choose these weighting measures independently. The measures $\hat{\pi}^{*}_{y_{i}}$’s are inherently determined by the invariant distribution $\pi_{\gamma^{*}}$ under the exploration policy $\gamma^{*}$. We first start with the following result which relates the approximation error to occupation measures.
@@
Under Assumption Assumption 2.2 and Assumption 3.1, given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}$:
$\left|J^{*}_{\beta}(x_{0})-\min_{v}\{Q^{*}(x_{0},v)\}\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\int_{B_{i}}\|x-x^{\prime}\|\mu_{\beta}(dx)\hat{\pi}_{y_{i}}(dx^{\prime}),$ (23)
@@
Assume that the state space $\mathds{X}\subseteq\mathbb{R}^{n}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
@@
where $V(x)=\|x\|_{1}^{m}$, $m>1$. Then, under Assumption 2.1 and Assumption 2.2, provided that the cost function $c$ is bounded, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathbb{X}$:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1/n(1-1/m)})(1-\beta)}\right),$ (26)
@@
$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
@@ 4 Concluding Remarks
For MDPs with unbounded state spaces we presented upper bounds on finite model approximation errors via optimizing the quantizers used for finite model approximations. We also considered implications on quantized Q-learning and the performance of policies obtained via Q-learning where the quantized state is treated as the state itself. We noted the distinctions between planning, where approximating MDPs can be independently designed, and learning, where approximating MDPs are restricted to be defined by invariant measures of Markov chains under exploration policies, leading to significant subtleties on quantizer design performance. Nonetheless, asymptotic near optimality can be established under both setups with explicit convergence rates.
@@ Appendix C Proof of Theorem 6
First consider the case with $n=1$, that is $\mathds{X}\subseteq\mathbb{R}$ Step 1: Partition $\mathds{X}$ into $M+1$ quantization bins $\{B_{1},B_{2},\ldots,B_{M},B_{M+1}\}$. The first $M$ bins cover a compact subset $\mathcal{K}=[-\frac{N}{2},\frac{N}{2}]\subset\mathds{X}$, and the last bin $B_{M+1}$ is an overflow bin that captures the rest of the state space outside $\mathcal{K}$. We apply a uniform quantizer to the compact region $\mathcal{K}$, dividing it into $M$ bins of equal length. The quantization width of each bin is:
@@
where $\mu_{\beta}$ is the normalized discounted occupation measure under a stationary policy $\gamma_{s}$. Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
@@
Step 3: In the overflow bin $B_{M+1}$, the state space may be unbounded, we take that the overflow bin is always mapped to state $x=0$: Thus,
@@
where $\hat{\pi}^{*}_{M+1}(dx^{\prime})$ is the normalized weighing measure over the bin $B_{M+1}$ and where the last step follows since we map the overflow bin directly to 0, i.e. $x^{\prime}=0$ with probability 1. Using Hölder’s inequality to bound the expected loss over $B_{M+1}$:
@@
Step 4: In the following, we bound the moment of the probability measure $\mu_{\beta}^{\gamma_{s}}$. Define the process $\{M_{t}\}$ for $t\geq 0$:
@@
with $M_{0}=V(x_{0})-b=|x_{0}|^{m}-b.$ Observe the following with respect to the filtration $\{\mathcal{F}_{t}\}$, where $\mathcal{F}_{t}=\sigma(X_{1},X_{2},\dots,X_{t})$ is the natural filtration:
@@
$\leq\frac{(1-\alpha)V(x_{t})}{(1-\alpha)^{t+1}}-\frac{b}{(1-\alpha)^{t+1}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\frac{b}{(1-\alpha)^{t+1}}=0,$
@@
$=(V(X_{0})-b)(1-\alpha)^{t}+b\sum_{k=0}^{t}(1-\alpha)^{k}=(V(X_{0})-b)(1-\alpha)^{t}+b\cdot\frac{1-(1-\alpha)^{t+1}}{\alpha}.$
@@
where $C$ is defined in (34). Combining the bounds from (31), (C) and (35):
@@
This completes the case with $n=1$. We now consider the case where $\mathds{X}\subseteq\mathbb{R}^{n}$ with $n\geq 1$.
Now, fix a side length parameter $N>0$. Choose an integer $k\geq 1$ and set
@@
which is the set of states outside the compact (granular) grid. We decompose the expected loss into two parts as in the proof of Theorem 6:
@@
Inside the overflow bin $B_{M+1}\subseteq\mathds{X}$, we take again that the overflow bin is directly mapped to state $x=0$ and thus,
@@
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\bigl(\|x\|_{1}+\|x^{\prime}\|_{1}\bigr)\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]$
@@
Repeating the arguments in Step 4,
@@
We select $N=2\bigl(C\,k\bigr)^{1/m}.$ By Markov’s Inequality:
@@
$\leq\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]}{(\tfrac{N}{2})^{m}}=\frac{C}{C\,k}=\frac{1}{k}.$
@@
$\Delta=\frac{N}{k}=2C^{\frac{1}{m}}k^{(1-\frac{1}{m})}.$
Inserting these bounds into (42):
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq 2nC^{1/m}\left(\frac{1}{k}\right)^{1-\frac{1}{m}}+C^{1/m}\left(\frac{1}{k}\right)^{1-\frac{1}{m}}\leq(2n+1)C^{1/m}k^{-(1-\frac{1}{m})}$
As $M\to\infty$, and so as $k\to\infty$, we have $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\to 0$. Thus,
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{\frac{1}{n}(1-\frac{1}{m})})(1-\beta)},$ (43)
where $n$ is the dimension of the state space, $M$ is the number of quantization bins, $m$ is the moment in the Lyapunov function, and $C$ is (defined in (34)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$. ∎
Appendix D Proof of Theorem 7
@@
Letting $\alpha:=(1-\mu(\mathds{X}))$, and repeating this step, we write $\|h_{k}\|_{\infty}\leq\|c\|_{\infty}\sum_{t=0}^{k-1}\alpha^{t}$. Finally, using the fact that the operator $\hat{T}$ is a contraction under the supremum norm under Assumption 2.3 with the fixed point $\hat{h}$, we conclude $\|\hat{h}\|_{\infty}\leq\frac{\|c\|_{\infty}}{1-\alpha}=\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$. Identical steps can be used to show that $\|h\|_{\infty}\leq\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$. Combining what we have so far, we write
$|j^{*}-\hat{j}|=\left|\int h(x)\mu(dx)-\int\hat{h}(x)\mu(dx)\right|\leq\int V(x)\mu(dx)\leq\left(\alpha_{c}+\frac{\|c\|_{\infty}\alpha_{T}}{\mu(\mathds{X})}\right)\int L(x)\pi_{\gamma_{s}}(dx).$
Appendix E Proof of Theorem 8
Proof.
Choose a collection of quantization bin $\{B_{i}\}_{i}^{M+1}$ such that the first $M$ bins quantize the compact set $(bk)^{1/m},(bk)^{1/m}]^{n}$ uniformly where $M=k^{n}$ and the last bin $B_{M+1}$ is the overflow bin that captures the rest of the state space. Then, the total expected distortion becomes:
As earlier, we choose $y_{M+1}=0$ (or any fixed point), so that $L(x)=\|x\|_{1}$ for $x\in B_{M+1}$. By our assumption, there exists a Lyapunov function $V$ that satisfies inequality (11). Thus, by [33, Theorem 4.2.5] (which builds critically on the Comparison Theorem [18, Theorem 14.2.2]), under any invariant probability measure $\pi$: $\int_{\mathds{X}}\|x\|_{1}^{m}\pi(dx)\leq b$
By following the same procedure as in the proof of Theorem 6, we obtain the following bound:
$\leq\Delta\ +\frac{b^{1/m}}{k^{1-1/m}}\leq\frac{2nb^{1/m}}{k^{1-1/m}}+\frac{b^{1/m}}{k^{1-1/m}}=\frac{(2n+1)b^{1/m}}{k^{1-1/m}}.$ (48)
$\left|\hat{J}_{avg}(x_{0})-J^{*}_{avg}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{(2n+1)b^{1/m}}{(M^{1/n(1-1/m)})},$ (49)
@@ Appendix F Proof of Theorem 11
Step1: We partition the state space similar to the quantization in Theorem 6 by choosing the compact set $\mathcal{K}=[-(Ck)^{1/m},(Ck)^{1/m}]$ with $k^{n}=M$, ($k=M^{1/n}$). Then, for any state $x\in\mathcal{K}$, the quantization error $L(x)$ satisfies:
$L(x)\leq{\Delta}=\frac{2n(Ck)^{1/m}}{k}=\frac{2nC^{1/m}}{k^{1-1/m}}=\frac{2nC^{1/m}}{M^{1/n(1-1/m)}}.$ (50)
@@
$\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}(\mathcal{K})=\frac{2nC^{1/m}}{M^{1/n(1-1/m)}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (52)
@@
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}\|x-x^{\prime}\|_{1}\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (53)
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\left(\|x\|_{1}+\|x^{\prime}\|_{1}\right)\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (54)
$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)]+\mathbb{E}_{\hat{\pi}_{M+1}}[\|X^{\prime}\|_{1}]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1}),$ (55)
@@
Observe that if $x_{0}\sim\pi_{\gamma^{*}}$ then $\mu_{\beta}^{\gamma_{s}}=\pi_{\gamma^{*}}$ and therefore the terms in the summation above will be identical, that is,
$\mathbb{E}_{\hat{\pi}^{*}_{M+1}}[\|X^{\prime}\|_{1}]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mu_{\beta}^{\gamma_{s}}(B_{M+1})\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)].$
@@
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|\mathds{1}_{B_{M+1}}(X)]=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\mathbb{E}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)|X_{0}=x]]\leq\hat{C}^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$
@@
$\hat{C}:=\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|x_{0}\|_{1}]^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$